OpenStudy (anonymous):

The population of an endangered species is decreasing so that 90% of the population remains after the passage of one year. If there are 320 individuals at the beginning of a study, then the function P = 320 (0.9)t gives the population at any time,t. (a) Make a table showing the population at the end of years 1,2,3 and 4. i got 288, 259,233,210 (b) In how many years will the population be reduced to 100 individuals? i set up the equation as 100=320(.9)^t i know that the answer is 11 but i can't seem to get the right answer.

OpenStudy (lgbasallote):

for (a) you should round off 259 to 260 and 233 to 234...after all there's no "half" animal for (b) \[100 = 320 (0.9)^t\] \[\frac{100}{320} = 0.9^t\] \[\ln (\frac{100}{320}) = t\ln (0.9)\] \[\frac{\ln (\frac{100}{320})}{\ln (0.9)} = t\] that should give you 11

OpenStudy (anonymous):

thank you my mistake was not doing the ln

OpenStudy (lgbasallote):

ahh i see. hope you learned your mistake :)